This is one question and question "7" need the information in "5" Thank you.

5 Cournot Suppose there are two departments selling economics degrees in one market competing fol- lowing the rules of the Cournot Oligopoly Model econ and man. econ. Suppose market demand for an economics degree is 7200-2p. Suppose both departments marginal cost is $3000 per degree. What is each department's residual demand curve? 5.1 5.2 What is each department's best response functions? 5.3 What is the Nash-Cournot equilibrium in this market? What is the price at this equilibrium? 5.4 What are profits for the two departments at this equilibrium? 5.5 Graph the marginal cost, demand, residual demand, and marginal revnue curve for the econ department. Show the equilibrium price and quantity for the econ department. 5.6
7 Bertrand Now suppose that instead of competing on quantities, the two departments are competing by setting prices following the Bertrand Oligopoly Model. Marginal cost remains the same and total market demand for economics degrees is 72 U. What is the Nash-Bertrand equilibrium for this market? 7.1 Now suppose that the two departments enter the market for online degrees with differentiated products. The marginal cost of each degree drops to S400 for each department. Suppose econ faces a demand function of ge = 10000-200Pe + 100pme . Suppose man. econ faces a demand function of qme 10000 200pme 100pe What is the Nash-Bertrand equilibrium now? 7.2
Answer #1

Profit of econ is maximized at

Profit (e) = (10000-200pe+100pme-400)pe

FOC: 9600-400pe+100pme=0

Pe= 9600+100pme / 400

Profit of man. econ. is maximized as

Profit( me) = (10000-200pme+100pe-400)pme

FOC: 9600-400pme+100pe =0

Pme= 9600+100pe / 400

Nash equilibrium will be

400pe = 9600+ 100(9600+100pe)/400

400pe = 9600 + 2400 + 25pe

375pe = 12000

Pe= 32

Pme= 32